OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1). You must show all work.

OpenStudy (anonymous):

do you know how to find slope?

OpenStudy (anonymous):

\[m = \frac{ y_2 - y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

standard form:\[y-y_1 = m(x-x_1)\]

OpenStudy (anonymous):

ya i know that.. but ive been working on this for an hour now and i need the answer!

OpenStudy (anonymous):

this problem should not take an hour to solve. simply plug in the points that you are given into the equation

OpenStudy (anonymous):

my test is due by 11

OpenStudy (anonymous):

i do not condone cheating on tests

OpenStudy (anonymous):

its homework my bad

OpenStudy (anonymous):

ha yeah

OpenStudy (anonymous):

seriously though, just plug in the points that were given to you

OpenStudy (anonymous):

ok do i plug them in for the standard form or the other one?

OpenStudy (anonymous):

well you have to find what the slope (m) is first then plug 1 set of points into the standard form

OpenStudy (anonymous):

or both?